1. The table below shows data on a subsample of participants in the 7th examination of the Framingham Offspring Study.

Characteristic

n

Sample Mean

Population Standard Deviation

Systolic Blood Pressure

10

121.2

11.1

Diastolic Blood Pressure

10

71.3

7.2

Total Serum Cholesterol

10

202.3

37.7

Weight

10

176.0

33.0

Height

10

67.175

4.205

Body Mass Index

10

27.26

3.10

Which of the following is the approximate margin of error for the 95% confidence level for Diastolic Blood Pressure

a. 0.03

b. 0.14

c. 4.5

d. 5.9
Answer #1

sample mean, xbar = 71.3
sample standard deviation, σ = 7.2
sample size, n = 10


Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, Zc = Z(α/2) = 1.96


ME = zc * σ/sqrt(n)
ME = 1.96 * 7.2/sqrt(10)
ME = 4.5

Option c) is answer